At the beginning of 20x1, Togo, Inc. entered into a finance lease to acquire equipment. The lease requires four annual payments of $25,663 beginning on December 31, 20x1. The present value of the lease payments, discounted at 8%, is $85,000. The leased asset is expected to be worthless at the end of the lease and Togo uses the straight-line depreciation method.

11. Based on the information above, Togo’s interest expense for the year ended 20x1 is closest to:

0

4,900

6,800


12. Based on the information above, Togo’s lease liability at the end of 20x2 is closest to:

0

45,000

66,000


13. Based on the information above, Togo’s depreciation expense for the year ended 20x1 is closest to:

0

21,000

28,000
Answer #1

11)correct option is "C" -6800

Interest payment =Lease liability carrying value * interest rate

                 = 85000*8%

                 = 6800

12)Correct option is "B" -45000

Date Payment Interest Principal repayment carrying value at end
1/1/20x1 85000
31/12/20x1 25663 6800 25663-6800= 18863 85000-18863= 66137
31/12/20x2 25663 66137*8%= 5290.96 25663-5290.96= 20372.04 66137-20372.04= 45764.96 (Closest to 45000)

13)

correct option is "B"-21000

Depreciation expense =Right of use asset /useful life

                 = 85000 /4

                  = 21250   (closest to 21000)